# A Scalable Bayesian Spatiotemporal Model for Water Level Predictions using a Nearest Neighbor Gaussian Process Approach

### v1

@@ Abstract
Obtaining accurate water level predictions are essential for water resource management and implementing flood mitigation strategies. Several data-driven models can be found in the literature. However, there has been limited research with regard to addressing the challenges posed by large spatio-temporally referenced hydrological datasets, in particular, the challenges of maintaining predictive performance and uncertainty quantification. Gaussian Processes (GPs) are commonly used to capture complex space-time interactions. However, GPs are computationally expensive and suffer from poor scaling as the number of locations increases due to required covariance matrix inversions. To overcome the computational bottleneck, the Nearest Neighbor Gaussian Process (NNGP) introduces a sparse precision matrix providing scalability without having to make inferential compromises. In this work we introduce an innovative model in the hydrology field, specifically designed to handle large datasets consisting of a large number of spatial points across multiple hydrological basins, with daily observations over an extended period. In this work, we investigate the application of a Bayesian spatiotemporal NNGP model to a rich dataset of daily water levels of rivers located in Ireland. The dataset comprises a network of 301 stations situated in various basins across Ireland, measured over a period of 90 days. The proposed approach allows for prediction of water levels at future time points, as well as the prediction of water levels at unobserved locations through spatial interpolation, while maintaining the benefits of the Bayesian approach, such as uncertainty propagation and quantification. Our findings demonstrate that the proposed model outperforms competing approaches in terms of accuracy and precision. When compared to single-basin approaches in the field of hydrology, our model achieves scalability and the ability to extend predictions across a broader range of river basins with minimal trade-off in predictive performance.
@@ 1 Introduction
In Ireland, rivers serve as a primary source of freshwater, providing drinking water to households, supporting industrial processes, and sustaining agriculture. The monitoring of water levels enables authorities to predict and manage water availability, especially during dry periods, as low levels can lead to water shortages, while high levels may require preventative measures to avert flooding and contamination in terms of sewage. Accurate predictions of river water levels are essential for effective water resource management and flood mitigation strategies.
@@
Ireland’s climate contributes to significant fluctuations in water levels with heavy precipitation in autumn and winter particularly. At these times there is heightened flood risk in many areas. Flood events lead to both direct economic losses, such as damage to infrastructure and property, and indirect losses through disruptions to economic activity. Between 1996 and 2015, the global economic losses from flooding were estimated at $597$ billion USD (Willner et al.,, 2018). Accurate water level prediction is therefore increasingly important in Ireland, given the rising demand for water across industrial and domestic sectors and the growing frequency of flood events (Charlton et al.,, 2006). Additionally, the intensification of precipitation patterns as a consequence of climate change increases the risks of flooding (Murphy et al.,, 2023). This highlights the critical need for accurate inputs to flood mapping models in order to protect communities and support resilient infrastructure planning.
@@
In the hydrology field, water level recordings have been traditionally collected through physical sensors strategically placed along the river stream, creating a spatial network of monitoring stations that collect data over time. The main challenge of making water level predictions is capturing the intrinsic spatiotemporal dynamics of river systems while accounting for related influencing factors. For instance, precipitation events increase water levels while evaporation has the opposite effect, contributing to water level reductions. Indeed, empirical studies have demonstrated that precipitation in the preceding day is the most significant factor for water level increases as precipitation makes its way through the soil or across non absorptive surfaces (Kenda et al.,, 2020; Ahmed et al.,, 2022). Another challenge inherent to the task of predicting water levels is the quality of the data. Physical sensors are prone to malfunctions, which can result in missing or unreliable measurements that introduce biases into the data. This has the potential to compromise predictive model accuracy and introduces additional complexity to the prediction task, since data-driven models often require effective pre-processing to handle such issues prior to being able to generate predictions.
@@
In the literature, two main approaches are employed for predicting water levels: hydrological simulations and/or data-driven models. The former uses physical models to simulate the water level dynamics through several obligatory inputs, such as soil and land-use related parameters, leaf area index and snow melt and frost parameters. However, these methods are often computationally costly, especially for large simulations and rely on assumptions of the input parameters derived from published literature (van der Knijff et al.,, 2010). For the latter, many data-driven methods have been applied to predict water levels within a space and time context using machine learning and statistical techniques. Pan et al., (2020) applied a convolutional neural network to predict daily water levels over an extended period (30 years) but only considered four water stations. More recently, Pagendam et al., (2023) tackled a large dataset consisting of weekly observations of groundwater depth using a deep neural network with two sub-architectures highlighting the advances of computational efficient methods. However, these machine learning models are less transparent, more difficult to interpret than statistical models and uncertainty quantification for model predictions are challenging.
@@
Space-time kriging has been widely applied to predict water levels under the frequentest (non-Bayesian) paradigm but can suffer from poor scaling as the number of locations increases (Varouchakis et al.,, 2022; Ruybal et al.,, 2019; Hoogland et al.,, 2010; Júnez-Ferreira and Herrera,, 2013; Kazemi et al.,, 2021). The Gaussian Process (GP) model is frequently used in the hydrological literature since it is capable of modeling complex space-time dependent covariance structures while providing a flexible approach for approximating smooth functions. Additionally, the approach has been widely studied under the frequentist and Bayesian paradigm. There is a strong connection between GPs and regression kriging, and these can be viewed as interchangeable in some scenarios. However, Varouchakis and Hristopulos, (2019) showed that GPs demonstrate superior performance when compared to kriging. In general, both GP and kriging approaches are susceptible to computational limitations when applied to large dataset due to the inversions of covariance matrices.
@@
The seminal paper of Vecchia, (1988) introduced an approximation for the GP that alleviates this computational limitation. As a result, the modified GP is scalable for larger datasets while keeping the benefits of a GP. Later, Datta et al., 2016a used this concept to propose the Nearest Neighbor Gaussian Process (NNGP) model and showed that this approximation is also a valid spatial process for the whole spatial domain without compromising inferential accuracy. More recently, the NNGP approach has achieved the state of the art in predictions at the second KAUST spatial statistics competition which compared different methodologies for spatial and spatiotemporal process (Abdulah et al.,, 2022). The application of NNGP approach to the field of climate science is a topic already addressed in the existing literature. For instance, (Datta et al., 2016b, ) investigated its predictive capabilities for a particulate matter air quality dataset. Mastrantonio et al., (2019) proposed a hierarchical multivariate NNGP model for predicting monthly precipitation, minimum and maximum temperatures. However, there is little to no consideration in the literature of its application to hydrological datasets.
@@
In the field of hydrology, where missing values are a common occurrence, non-Bayesian approaches, such as those discussed above, have the disadvantage of requiring the use of an auxiliary model to address the missingness of data. In contrast, Bayesian approaches offer a natural solution for imputation through the specification of prior distributions. Molinos et al., (2015) proposed a Bayesian harmonic regression model to predict monthly water levels in Irish lakes considering an extended period (1974 - 2012) while accounting for seasonal effects. While this study provided insights into long-term water level trends, it analysed recordings from only 28 spatial locations. Related to water levels, Santos-Fernandez et al., (2022) proposed various Bayesian vector autoregressive (VAR) models to predict daily water temperature in rivers, considering covariance functions depending on stream networks and Euclidean distances. However, despite the innovation introduced by the authors, computational issues arise again as the method scales poorly for a large number of spatial points.
@@
Despite the development of spatiotemporal models in the field of hydrology, there have been few successful studies addressing the challenge of predicting water levels across a dense network of monitoring stations with extensive temporal records, all while maintaining computational efficiency and model performance. This task remains complex, as it requires a model capable of simultaneously overcoming the computational limitations, capturing the space-time dependencies and accounting for prediction uncertainty and data gaps. In order to address this challenge, we propose a Bayesian spatiotemporal NNGP model to predict daily water levels considering a large number of stations located in Ireland recorded over an extended period. The dataset consists of a dense network of 301 stations each measured over 90 days. The model incorporates the past-day precipitation as a predictor with an additive spatiotemporal effect following a first-order autoregressive process. The proposed approach allows for predicting both monitored stations in a future day and predicting unobserved locations through the use of interpolation.
@@
The remainder of the paper is structured as follows. In Section 2 we describe the water level and precipitation dataset in Ireland. In Section 3 the NNGP definition and detailed explanation of the proposed spatiotemporal NNGP model is presented. In Section 4, we present the strategy and metrics used to assess model performance, followed by the results of applying the model to both synthetic and real-world water level data. Following that, in Section 5 we compare our proposed model with alternative approaches documented in the literature. In Section 6 the limitations of this study are discussed. Finally, the Section 7 concludes the paper with a summary of findings and suggestions for future research.
@@ 2 Water level and precipitation records in Ireland
The Office of Public Works (OPW) collects water level and additional variables in a daily frequency for an extensive hydrometric network of 380 monitoring stations located on rivers, lakes and tidal locations throughout the Republic of Ireland. The data is accessible via the Application Programming Interface (API) which can be accessed at https://waterlevel.ie/ (OPW,, 2021). For this study, we selected only those stations located on rivers and extracted the water level recordings (in meters) for the period between 03/01/2022 and 02/04/2022. To ensure the reliability of the records, stations with more than 30% of missing data were ignored, resulting in 301 stations. Given the substantial variability in the average water level across rivers, we subtracted the water level recordings of each station by their respective means, resulting in a zero-centered distribution (see Figure 1). The distances between locations were calculated using the great circle distance to account for the curvature of the earth, thus enhancing precision and ensuring a representation in kilometres.
@@
Figure 1: Zero-centered water levels (in meters) observed at stations across Ireland between 4 February 2022 and 7 February 2022.
@@
The Irish Meteorological Service, Met Éireann provides forecasting and access to historical data on several weather variables, including 24 hours precipitation accumulation, temperature, wind speed, soil moisture and potential evapotranspiration. We extracted the precipitation variable from the available weather stations, covering the same one-year period as the water level data. The data retrieval was facilitated using the API available at https://marie.ie/repo/met-eireann-api/ (Marie,, 2023).
@@
The observed precipitation did not align with the coordinates of river stations (see Figure 2). To address this spatial misalignment, the Inverse Distance Weighting (IDW) interpolation was applied to estimate the rainfall at each river station. The IDW is a deterministic interpolation method where the estimated values are calculated as weighted averages of observed values, with the weights being inversely proportional to the distance between each observed point and the new point. The methodology has been demonstrated to be effective for interpolating precipitation data, as evidenced by (Chen and Liu,, 2012).
@@
Figure 2: Map of Ireland illustrating the spatial misaligment between river (red) and hydroclimate stations (green).
In order to address the heavily right-skewed distribution of precipitation and strengthen its correlation with water level, we applied the square root transformation. Furthermore, in recognition of the delayed impact of precipitation on water level, we considered the precipitation from the previous day (lagged precipitation) as a predictor (see Figure 3). For clarity, we will refer to this transformed and lagged variable as precipitation throughout the analysis.
Figure 3: Left: precipitation before (square root and lag) transformation versus zero-centered water level. Right: precipitation after the transformation versus zero-centered water level.
The left panel in Figure 4 illustrates the temporal evolution of water level across all stations throughout the study period. The figure reveals a general pattern among the majority of the stations with peak records occurring in February. The accompanying spatiotemporal semivariogram (right panel in Figure 4) provides further support for the trend pattern observed among the stations. The semivariance rapidly increases with distance and then flattens out, indicating a "sill" where the spatial correlation diminishes. However, the semivariance persists consistently over time, suggesting that while spatial proximity significantly influences water levels until a distance is reached, the temporal dynamics exhibit a long-lasting effect. The stations following different patterns to this may be attributed to site-specific influences or unreliable measurements.
Figure 4: Left: temporal evolution of water level. Right: empirical spatiotemporal semivariogram for the considered monitoring stations.
When inspecting simultaneously the water level and precipitation records, we see high water levels in February as a result of intense precipitation (Figure 5). In Ireland, the wet season typically occurs in February and March. Despite March being classified as a wet month (the precipitation records at the beginning of the month can be comparable to February), the distribution of water levels resembles the observed levels in January. This suggests that factors beyond precipitation are likely influencing water level variations.
@@
Figure 5: Left: boxplot of water level above the mean level (meters). Right: precipitation (millimetres) by month.
@@
The impact of water level drivers beyond precipitation is clearly demonstrated when examining the fluctuations in water levels across stations over time (Figure 6). However, precipitation peaks consistently occur in late February and early March, water levels show significant variation in how they respond to these peaks across different stations. This pattern indicates that additional hydroclimate factors such as soil saturation and evapotranspiration rates may be shaping the observed water level dynamics alongside precipitation. Furthermore, stations with similar water level variations may be situated in close proximity to one another, indicating the presence of a spatial correlation (as previously highlighted in the right panel in Figure 4).
@@
Figure 6: Water level above the mean level (meters) and precipitation (millimetres) at six river monitoring stations.

### v2

@@ Abstract
Obtaining accurate water level predictions are essential for water resource management and implementing flood mitigation strategies. Several data-driven models can be found in the literature. However, there has been limited research with regard to addressing the challenges posed by large spatio-temporally referenced hydrological datasets, in particular, the challenges of maintaining predictive performance and uncertainty quantification. Gaussian Processes (GPs) are commonly used to capture complex space-time interactions. However, GPs are computationally expensive and suffer from poor scaling as the number of locations increases due to required covariance matrix inversions. To overcome the computational bottleneck, the Nearest Neighbor Gaussian Process (NNGP) introduces a sparse precision matrix providing scalability without compromising its inferential capabilities. In this work, we introduce an innovative model in the hydrology field, specifically designed to handle large datasets consisting of a large number of spatial points across multiple hydrological basins, with daily observations over an extended period. We investigate the application of a Bayesian spatiotemporal NNGP model to a rich dataset of daily water level of rivers located in Ireland. The dataset comprises a network of 301 monitoring stations situated in various basins across Ireland, measured over a period of 90 days. The proposed approach allows for prediction of water level at future time points, as well as the prediction at unobserved locations through spatial interpolation. Furthermore, the Bayesian approach provides the benefits in terms of uncertainty propagation and quantification, which are of considerable importance to the hydrology field. Our findings demonstrate that the proposed model outperforms competing scalable approaches in terms of accuracy and precision. Moreover, the proposed model enables predictions across broader range of water bodies with minimal trade-off in predictive performance.
@@ 1 Introduction
In Ireland, rivers serve as a primary source of freshwater, providing drinking water to households, supporting industrial processes, and sustaining agriculture. The monitoring of water level enables authorities to predict and manage water availability, especially during dry periods, as low levels can lead to water shortages, while high levels may require preventative measures to avert flooding and contamination in terms of sewage. Accurate predictions of water level are essential for effective water resource management and flood mitigation strategies.
@@
Ireland’s climate contributes significantly to water level fluctuations with heavy precipitation during autumn and winter being a major contributing factor. At these times there is heightened flood risk in many areas. Flood events lead to both direct economic losses, such as damage to infrastructure and property, and indirect losses through disruptions to economic activity. Between 1996 and 2015, the global economic losses from flooding were estimated at $597$ billion USD (Willner et al.,, 2018). Accurate water level prediction is therefore increasingly important in Ireland due to the rising frequency of flood events, as well as to the increasing demand for water across industrial and domestic sectors (Charlton et al.,, 2006). Additionally, the effects of climate change have intensified precipitation patterns, therefore increasing the risks of flooding (Murphy et al.,, 2023). This highlights the critical need for models to accurately predict water level in order to protect communities and support resilient infrastructure planning.
@@
In the hydrology field, water level recordings have been traditionally collected through physical sensors strategically placed along the river stream, creating a network of monitoring stations that collect data over time. The main challenge of making water level predictions is capturing the intrinsic spatiotemporal dynamics of river systems while accounting for related influencing factors. For instance, precipitation events increase water levels while evaporation has the opposite effect, contributing to water level reductions. Indeed, empirical studies have demonstrated that the precipitation in the preceding days are the most significant factor in water level increases (Kenda et al.,, 2020; Ahmed et al.,, 2022).
Another challenge inherent to the task of predicting water level is the quality of the data. Physical sensors are prone to malfunctions, which can result in missing or unreliable measurements that introduce biases into the data. This introduces additional complexity to the prediction task, since data-driven models often require effective pre-processing to handle such issues prior to being able to generate predictions. If not properly managed, poor data quality can significantly compromise the model’s predictive capabilities.
@@
In the literature, two main approaches are employed for predicting water level: hydrological simulations and/or data-driven models. The former uses physical models to simulate the water level dynamics through several obligatory inputs, such as soil and land-use related parameters, leaf area index and snow melt and frost parameters. However, these methods are often computationally costly, especially for large simulations and rely on assumptions of the input parameters derived from published literature (Knijff et al.,, 2010). For the latter, many data-driven methods have been applied to predict water level within a space and time context using machine learning and statistical techniques. Pan et al., (2020) applied a convolutional neural network to predict daily water level over an extended period (30 years) but only considered four water stations. More recently, Pagendam et al., (2023) tackled a large dataset consisting of weekly observations of groundwater depth using a deep neural network with two sub-architectures highlighting the advances of computational efficient methods. However, the mentioned machine learning approaches are limited to predictions only at monitoring sites and often lack the interpretability and transparency as opposed to statistical models. Furthermore, these methods typically do not provide rigorous quantification of predictive uncertainty.
@@
Space-time kriging has been widely applied to predict water level under the frequentest (non-Bayesian) framework (Varouchakis et al.,, 2022; Ruybal et al.,, 2019; Hoogland et al.,, 2010; Júnez-Ferreira and Herrera,, 2013; Kazemi et al.,, 2021). However, the approach suffer from poor scaling as the number of locations increases. The Gaussian Process (GP) model is frequently used in the hydrological literature since it is capable of modelling complex space-time dependent covariance structures while providing a flexible approach for approximating smooth functions. Additionally, the approach has been widely studied under the frequentist and Bayesian inference. There is a strong connection between GPs and regression kriging, and these can be viewed as interchangeable in some scenarios. However, Varouchakis and Hristopulos, (2019) showed that GPs demonstrate superior performance when compared to kriging. In general, both GP and kriging approaches are susceptible to computational limitations when applied to large dataset due to the inversions of covariance matrices.
@@
The seminal paper of Vecchia, (1988) introduced an approximation for the GP that alleviates its computational limitations, making it alternative and feasible approach for large datasets relying on GPs. Based on this concept, Datta et al., 2016a proposed the Nearest Neighbor Gaussian Process (NNGP) model, showing that this approximation is a valid spatial process over the whole spatial domain without compromising its inferential capabilities. More recently, the NNGP approach has been recognised as the state of the art in terms of prediction for spatial and spatiotemporal problems (Abdulah et al.,, 2022). The application of NNGP approach to the field of climate science is a topic already addressed in the existing literature. For instance, (Datta et al., 2016b, ) investigated its predictive power for a particulate matter air quality dataset. Mastrantonio et al., (2019) proposed a hierarchical multivariate NNGP model for predicting monthly precipitation, minimum and maximum temperatures. However, there is little to no consideration in the literature of its application to hydrological datasets.
@@
In the field of hydrology, where missing values are a common occurrence, non-Bayesian approaches, such as those discussed above, have the disadvantage of requiring the use of an auxiliary model to address the missingness of data. In contrast, Bayesian approaches offer a natural solution for imputation through the specification of prior distributions. Molinos et al., (2015) proposed a Bayesian harmonic regression model to predict monthly water level in Irish lakes considering an extended period (1974 - 2012) while accounting for seasonal effects. While this study provided insights into long-term water level trends, it analysed recordings from only 28 spatial locations. Related to water level prediction, Santos-Fernandez et al., (2022) proposed a Bayesian space-time model to predict daily water temperature in rivers considering covariance functions depending on stream and Euclidean distance. Despite the innovation introduced by the authors, computational issues arise again as the method rely on GPs. Moreover, the approach requires prior informations of the dendritic stream network obtained through an extensive processing of geospatial hydrological databases.
@@
Despite the development of spatiotemporal models in the field of hydrology, a few successful studies have successfully addressed the challenge of predicting water level across a dense network of monitoring stations with extensive temporal records. Such studies are even more scarce when considering results encompassing both high predictive performance and computational efficiency. This task remains challenging due to several factors. As it requires an approach capable of simultaneously overcoming the computational limitations, capturing the complex space-time dependencies, accounting for prediction uncertainty and handling data gaps. In order to address this challenge, we propose a Bayesian spatiotemporal NNGP model to predict daily water level considering a large number of stations located in Ireland recorded over an extended period. The dataset consists of a dense network of 301 stations measured over 90 days. The model describe the water level variability in terms of a single predictor (five-days moving average of precipitation) and an additive spatiotemporal effect following a first-order autoregressive process to capture the space-time process. The proposed approach allows for predicting both monitored stations in a future day and predicting unseen locations through spatial interpolation.
@@
The remainder of the paper is structured as follows. In Section 2 we describe the Irish water level and precipitation dataset. Section 3 presents the definition of NNGP, followed by the formulation and detailed explanation of our proposed spatiotemporal NNGP model. In Section 4 we outline the evaluation strategy and performance metrics. Model results are then presented for a synthetic dataset, followed by the application to the Irish water level data. In Section 5 we compare our proposed model with alternative approaches documented in the literature. In Section 6 the limitations of this study are discussed. Finally, Section 7 concludes the paper with a summary of findings and suggestions for future research.
@@ 2 Water level and precipitation records in Ireland
The Office of Public Works (OPW) collects water level in daily and sub-daily frequency for an extensive hydrometric network of 301 monitoring stations situated on rivers, lakes and tidal locations throughout the Republic of Ireland. The monitoring stations are primarily located in areas susceptible to flooding or considered critical for water resource management. For this study, we selected only those stations located on rivers and extracted the daily average water level recordings (in meters) for the period from 03/01/2022 to 02/04/2022. The water level measurements correspond to the daily average water depth above the zero level on the local staff gauge. The data was extracted via the Application Programming Interface (API) available at https://waterlevel.ie/ (OPW,, 2021). Given the substantial variability in the water level across rivers and to facilitate the model fitting, we centred the water level measurements by subtracting the mean value at each station. This transformed variable represents the water level relative to the mean stage, which we adopt as the response variable in the analysis. For simplicity, we refer to this variable as water level whenever the context is clear. The water level distribution across the monitoring stations for a sample of days is illustrated in Figure 1. The distances between stations were calculated using the great circle distance to account for the curvature of the earth, thus enhancing precision and ensuring a representation in kilometres.
@@
Figure 1: Water level relative to the mean stage (meters) observed at stations across Ireland between 4 February 2022 and 7 February 2022.
@@
The Irish Meteorological Service, Met Éireann provides forecasting and access to historical data on several weather variables, including 24 hours precipitation accumulation, temperature, wind speed, soil moisture and potential evapotranspiration. However, for this study, only precipitation data were public available for the relevant time frame. We extracted the precipitation variable for the available weather stations, covering the same period as the water level data. The data retrieval was facilitated using the API available at https://marie.ie/repo/met-eireann-api/ (Marie,, 2023).
@@
The observed precipitation did not align spatially with the coordinates of river monitoring stations (see Figure 2). To address this spatial misalignment, the Inverse Distance Weighting (IDW) interpolation was applied to estimate the precipitation at the location of each river station. The IDW is a deterministic interpolation method which the estimated values is computed based on the weighted averages of observed values, where the weights are inversely proportional to the distance between the spatial points. This methodology has been demonstrated to be effective for interpolating precipitation data, as evidenced by Chen and Liu, (2012).
@@
Figure 2: Map of Ireland illustrating the spatial misalignment between river (red) and hydroclimate stations (green).
In order to better capture the relationship between precipitation and water level, particularly with regard to the delayed hydrological response, a five-day moving average was applied to the daily precipitation data. This smoothing technique strengthens the linear correlation with water level fluctuations and helps to capture the rainfall peaks. The resulting smoothed precipitation series was then used as a covariate in the analysis (see Figure 3). For clarity, we refer to this smoothed variable as precipitation throughout the remainder of the manuscript.
Figure 3: Left: relationship between daily precipitation and water level relative to the mean stage. Right: strengthen relationship between precipitation trends based on five-day moving average and water level relative to the mean stage.
As demonstrated in Figure 4, the temporal evolution of water level across the monitoring stations during the study period is illustrated. A common pattern is evident across most stations, with pronounced peaks occurring in early and late February to early March. However, some stations exhibit deviations from this general trend, either with attenuated or delayed peaks.
Figure 4: Temporal evolution of water level for the monitoring stations.
A monthly analysis of water level and precipitation reveals elevated water level in February, primarily attributed to prolonged periods of heavy rainfall. In a similar way, heightened water level are evident in March, aligning with notable precipitation peaks as indicated by the substantial number of outliers observed in the precipitation data (see Figure 5). These precipitation patterns are consistent with the climatology of Ireland, where the wet season typically occurs during the winter months and gradually diminishes as it transitions into spring, contributing to a subsequent decline in water levels.
@@
Figure 5: Distribution of precipitation and water level relative to the mean stage by month.
@@
The impact of water level drivers beyond precipitation is clearly demonstrated when examining the fluctuations in water level across stations over time (Figure 6). Although all six stations experienced peaks in precipitation during late February and early March, the water level response to these events varied notably across stations. This pattern indicates that additional hydroclimate factors such as soil saturation and temperature may be affecting the water level dynamics alongside precipitation.
@@
Figure 6: Water level relative to the mean stage (meters) and precipitation (millimetres) at six river monitoring stations.